OpenStudy (anonymous):

Simplify by combining like terms. 9 times g plus thirteen times g minus 4 minus sixteen times g minus 2 times g 0 nineteen times g minus twenty 4 times g minus 4 g

OpenStudy (anonymous):

will you please type out your problem

OpenStudy (anonymous):

\[9g+13g-4-16g-2g\]

OpenStudy (anonymous):

Combine and simplify @saundra789

OpenStudy (anonymous):

You have to combine like terms, group all the similar terms together, giving you 9g+13g-16g-2g-4. Now, combine all the g's

OpenStudy (anonymous):

\[9+13-16-2=22-16-2=8-2=6\]

OpenStudy (anonymous):

@satellite73 it's 6-2=4 at the last sentence